OpenStudy (anonymous):

During its first week of business, a market sold a total of 108 apples and oranges. The second week, five times the number of apples and three times the number of oranges were sold. A total of 452 apples and oranges were sold during the second week. Determine how many apples and how many oranges were sold the first week.

OpenStudy (taramgrant0543664):

So you can start by making two equations. We will let A represent the number of apples and R represent the number of oranges. The first week we had a total of 108 apples and oranges sell so the first equation can look like: A+R=108 The second week we had five times the apples and three times the oranges for a total of 452 so our equation can look like this: 5A+3R= 452 To solve we can rearrange the first equation to A=108-R and put that into the second equation so: 5(108-R)+3R=452 540-5R+3R=452 -2R=-88 R=44 We can then put the number of oranges into the first equation to determine the number of apples: A+R=108 A+44=108 A=64